# ART: Adaptive Resampling-based Training for Imbalanced Classification

### v1

@@
Abstract.
Traditional resampling methods for addressing class imbalance in supervised classification typically use fixed sampling distributions, either uniformly undersampling the majority class or oversampling the minority class. These static strategies fail to account for changes in class-wise learning difficulty during the training process, which can limit the overall performance of the model.
This paper proposes an Adaptive Resampling-based Training (ART) method that periodically updates the distribution of the training data based on the class-wise performance of the model. Specifically, ART uses class-wise macro F1 scores, computed at fixed intervals, to determine the degree of resampling to be performed.
@@ Abstract.
Experimental results across diverse class-imbalanced benchmark datasets like the Pima Indians Diabetes dataset and the Yeast dataset demonstrate that ART consistently outperforms both resampling-based and algorithm-level methods, including Synthetic Minority Oversampling Technique (SMOTE), NearMiss Undersampling, and Cost-sensitive Learning on binary as well as multi-class classification tasks with varying degrees of imbalance.
@@
In most settings, these improvements are statistically significant. On tabular datasets, gains are significant under both paired t-tests and Wilcoxon signed-rank tests (p ¡ 0.05), while performance on text and image tasks remains consistently favorable. Compared to training on the original imbalanced data, ART improves macro F1 by an average of 2.64 percentage points across all tested tabular datasets. Unlike existing methods, whose performance varies by task, ART consistently delivers the strongest macro F1 score, making it a reliable and broadly effective choice for imbalanced classification problems.
@@
Adaptive Resampling-based Training, Difficulty-aware Learning, Class Imbalance, Resampling Strategies, Supervised Classification, Data Mining, Machine Learning
copyright: noneccs: Computing methodologies Batch learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
The highly challenging task of training classifiers on imbalanced datasets is a long-standing problem in supervised learning. In real-world applications, like fraud detection, medical diagnosis, and rare event prediction, the class distributions are often highly skewed, leading standard learning algorithms to be heavily biased towards the majority class. To mitigate this problem, many techniques have been proposed. Existing approaches include resampling-based methods, which either resample statically before the training process or dynamically while training. Static approaches like Random Over-Sampling (ROS) (Chen et al., 2024), Random Under-Sampling (RUS) (Chen et al., 2024), Synthetic Minority Oversampling Technique (SMOTE) (Chawla et al., 2002), and nearmiss undersampling (Mani, 2003) fail to consider the changing difficulty of identifying a class as training progresses. Current dynamic approaches focus on instance-level hardness and often fail to account for imbalance in data. Other approaches are algorithmic, such as cost-sensitive learning (Elkan, 2001a), Online Hard Example Mining (OHEM) (Shrivastava et al., 2016), and focal loss (Lin et al., 2018), which modify the loss function to assign higher costs to misclassifying or minority class examples, thereby increasing the gradient contribution from these instances during optimization. More importantly, these methods fail to consider the changing difficulty of identifying a class as training progresses. While being effective in some cases, all of these methods share common assumptions, as they either (i) treat imbalance as a static property, fixed before the training begins, (ii) focus on instance-level hardness rather than class-level, or (iii) rely on complex heuristics to assign class-specific loss weights. These assumptions ignore the fact that class difficulty evolves during learning (Sinha et al., 2022; Toneva et al., 2019). A class that initially seems difficult due to low representation may, over time, become easier to learn as the model adapts. Static correction cannot adapt to these dynamics, and often leads to overfitting on the minority class or wasting capacity by oversampling far beyond what is actually needed.
To this end, we propose ART, a resampling-based method that uses the class-wise macro F1-score as a simple heuristic for adjusting the data distribution. To the best of our knowledge, this is a novel approach that dynamically resamples data based on model performance. ART periodically monitors class-wise performance and updates the sampling distribution to emphasize underperforming classes. This class-level focus avoids the instability associated with instance-level difficulty estimation (Sinha et al., 2022) and treats class difficulty as dynamic rather than static throughout training. Figure 1 shows how the sampling distribution of classes changes based on the macro F1-score as training progresses. Using a class-wise metric as a resampling heuristic offers a simpler and more robust alternative to complex fixed heuristics. Figure  2 shows how training progresses under ART. Unlike  (Sinha et al., 2022), which uses both resampling and a modified loss function, ART demonstrates that performance metric-based resampling alone outperforms existing methods. Also, ART utilizes a balance of performance metric-based sampling and the original dataset priors rather than completely ignoring the original data distribution. We demonstrate that ART consistently outperforms existing methods across classification tasks, achieving superior accuracy, recall, precision, and F1-scores. We also show that ART performs better than the baseline at varying levels of imbalance. Furthermore, we evaluate ART on classification tasks across tabular, image, and text datasets to demonstrate its modality-agnostic nature.
Figure 1. Adaptive resampling in ART is based on class-wise macro F1-score. The F1-score is used to estimate class difficulty, and the data distribution is updated every bfbf epochs to emphasize underperforming classes.
@@ 1. Introduction
The image shows the number of samples of the majority and minority class using red and green colours. It shows the initial distribution of the dataset and the distribution after resampling. It shows how resampling in ART is based on the F1-scores and that resampling increases the number of samples of the underperforming class in the training data
2. Related Work
@@
Resampling-based methods: A common approach to handle class imbalance is to rebalance the data distribution before or during training. Classic techniques include ROS, which simply duplicates minority-class examples, and RUS, which deletes majority-class examples. ROS and RUS are simple resampling methods and have been shown to be surprisingly competitive on some tasks. More sophisticated oversampling methods generate synthetic examples. These include methods like SMOTE (Chawla et al., 2002), which interpolates new minority examples in the feature space. Many SMOTE variants aim to focus sampling on hard or informative regions. For instance, Borderline-SMOTE (Han et al., 2005) generates points near class boundaries, and ADASYN (He et al., 2008) adjusts sampling density based on local minority density. There are also some recent methods like Safe-Level-SMOTE (Bunkhumpornpat et al., 2009) and k-means–SMOTE (Douzas et al., 2018) that combine clustering or noise filtering with synthetic generation. In practice, oversampling can introduce noisy or redundant points, so hybrid methods that combine oversampling with cleaning have been proposed in (Batista et al., 2004a). For example, SMOTE-ENN and SMOTE-Tomek, as described in (Batista et al., 2004b), use Edited Nearest Neighbors or Tomek-link removal to prune ambiguous samples after SMOTE. It has been shown that these hybrid methods yield well-defined class clusters and often outperform vanilla undersampling, especially when the minority class is minimal (Batista et al., 2004a, b).
@@
Loss-based methods: Rather than altering the data distribution, many techniques modify the learning algorithm or loss function. Classic cost-sensitive learning assigns a higher weight to the loss component of minority classes (Elkan, 2001b). One can weigh the loss inversely by class frequency, or use a pre-specified cost matrix (Elkan, 2001a). However, determining optimal costs can be difficult for complex models. Focal loss (Lin et al., 2018) addresses class imbalance in deep learning by changing the loss contribution of easy (well-classified) examples, thereby encouraging the model to focus on harder instances. Similarly, OHEM (Shrivastava et al., 2016) dynamically selects the examples with the highest loss within each mini-batch and focuses training on those “hard” samples rather than the full dataset. Although OHEM originated in computer vision, the idea of mining difficult instances can be applied in any domain.
@@
Techniques like self-paced learning (Kumar et al., 2010) or focal loss adjust selection based on how confidently the model predicts each sample. Some recent works, like AdaSampling frameworks (Zhang et al., 2019), iteratively update example weights by model feedback, although many such methods focus on label-noise scenarios. Margin-based sampling methods, like MSYN (Fan et al., 2011), explicitly choose instances near class boundaries to improve learning when the classifier has small margins on some samples. In general, instance-level adaptive sampling contrasts with class-level rebalancing because it can upweight or downweight individual examples even within the same class.
Multi-modal Imbalanced Learning: In practice, imbalanced methods are applied across domains (vision, NLP, tabular data, etc.). Techniques like random sampling and SMOTE assume only vector inputs and have been used extensively in tabular and image datasets. Modern deep-learning loss reweighting (e.g., LDAM+DRW, focal loss) was popularized on vision tasks but extends naturally to NLP or other modalities. For instance, focal loss has been applied in text classification to handle rare classes, and cost-sensitive learning is common in medical diagnosis or fraud detection on tabular data. Class-balanced sampling and ensemble methods like Balanced Meta-Softmax (Ren et al., 2020) and LDAM+DRW have been evaluated on both image benchmarks and real-world text or tabular tasks. Moreover, some recent works explicitly benchmark multiple modalities: for example, Cui et al. (Cui et al., 2019) and Cao et al. (Cao et al., 2019) validated their class-balanced and margin-based losses on both image recognition and text classification tasks.
Figure 2. Overview of the training procedure in ART. The process begins with uniform initialization of class weights. Training proceeds in epochs, with periodic updates to class weights every bf epochs based on per-class macro F1-scores evaluated on a validation set.
[…]
Flow chart showing the training in ART. Training start by initialising class weights equally for each class and storing the dataset priors. We perform training for 1 epoch. After each epoch, check if early stopping has been reached. If yes, then stop training. Else, check if it has been bf epochs since the last resampling. If yes, then calculate new class weights using the macro F1-score, resample the data and continue training.
3. Methodology
@@
See Algorithm 1 for how these initial weights are blended with the empirical priors via convex interpolation to form the first sampling distribution.
The classification model used in this study is a parameterized function, defined by:
@@
3.1. Overview
The main aim of ART is to enhance model performance in classification tasks by explicitly addressing variations in predictive performance across classes. ART accomplishes this by utilizing an adaptive sampling mechanism informed directly by the current performance of the classifier, rather than relying solely on fixed empirical class frequencies. In contrast to conventional static resampling methods, ART periodically revises the sampling distribution by assessing and quantifying the classification difficulty of individual classes based on validation metrics. The algorithm for training and resampling under ART is explained in Algorithm  1. Classes identified as challenging, indicated by lower predictive performance, receive increased sampling priority. Consequently, ART systematically reallocates training resources towards improving performance on classes exhibiting greater classification complexity.
@@ 3.1. Overview
  1. (1)
    the blending constant, denoted by $c\in[0,1]$, which controls the balance between sampling based on empirical class priors and the adaptive distribution derived from current class performance metrics;
  2. (2)
    the boosting frequency, denoted by $bf\in\mathbb{Z}^{+}$, which specifies the interval, measured in epochs, at which the sampling distribution is updated.
These hyperparameters influence how ART adjusts its sampling strategy over the course of training, and affects both the frequency and intensity of its response to class-level learning difficulty changes.
@@
3.2. Performance-based Sampling
@@
These computed difficulty scores reflect the inverse performance relationship and hence, emphasize classes with poor predictive performance. To create a valid probability distribution from these difficulty scores, we normalize them across all classes, resulting in a performance-based sampling weights given by:
@@
3.3. Blending with Class Priors

### v2

# ART: Adaptive Resampling-based Training for Difficulty-aware Classification

@@
Abstract
Traditional resampling methods for addressing class imbalance in supervised classification typically use fixed sampling distributions, either uniformly undersampling the majority class or oversampling the minority class. These static strategies fail to account for changes in class-wise learning difficulty during the training process. This paper proposes an Adaptive Resampling-based Training (ART) method that periodically updates the distribution of the training data based on the model’s class-wise performance. Specifically, ART uses class-wise macro F1 scores computed at fixed intervals to determine the degree of resampling to perform.
@@ Abstract
Experimental results across diverse class-imbalanced benchmark datasets demonstrate that ART consistently outperforms both resampling-based and algorithm-level methods, including Synthetic Minority Oversampling Technique, nearmiss undersampling, and cost-sensitive learning on binary as well as multi-class classification tasks with varying degrees of imbalance.
@@
In most settings, these improvements are statistically significant. On tabular datasets, gains are significant under both paired t-tests and Wilcoxon signed-rank tests (p ¡ 0.05), while performance on text and image tasks remains consistently favorable. ART improves macro F1 by an average of 2.64 percentage points across all tested tabular datasets. Unlike existing methods, ART consistently delivers the highest macro F1 score, making it a reliable and broadly effective choice for imbalanced classification problems.
keywords:
@@
Adaptive Resampling-based Training , Difficulty-aware Learning , Class Imbalance , Resampling Strategies , Supervised Classification , Data Mining
journal: Machine Learning with Applications
@@
[1]organization=School of Computer Science and Engineering, Vellore Institute of Technology,addressline=Katpadi-Tiruvalam Road, city=Vellore, postcode=632014, state=Tamil Nadu,country=India
1 Introduction
The highly challenging task of training classifiers on imbalanced datasets is a long-standing problem in supervised learning. In real-world applications, like fraud detection, medical diagnosis, and rare event prediction, the class distributions are often highly skewed, leading standard learning algorithms to be heavily biased towards the majority class. To mitigate this problem, many techniques have been proposed. Existing approaches include resampling-based methods, which either resample statically before the training process or dynamically while training. Static approaches like Random Over-Sampling (ROS) (Chen et al., 2024), Random Under-Sampling (RUS) (Chen et al., 2024), Synthetic Minority Oversampling Technique (SMOTE) (Chawla et al., 2002), and nearmiss undersampling (Mani, 2003) fail to consider the changing difficulty of identifying a class as training progresses. Current dynamic approaches focus on instance-level hardness and often fail to account for imbalance in data. Other approaches are algorithmic, such as cost-sensitive learning (Elkan, 2001b), Online Hard Example Mining (OHEM) (Shrivastava et al., 2016), and focal loss (Lin et al., 2018), which modify the loss function to assign higher costs to misclassifying or minority class examples, thereby increasing the gradient contribution from these instances during optimization. More importantly, these methods fail to consider the changing difficulty of identifying a class as training progresses. While being effective in some cases, all of these methods share common assumptions, as they either (i) treat imbalance as a static property, fixed before the training begins, (ii) focus on instance-level hardness rather than class-level, or (iii) rely on complex heuristics to assign class-specific loss weights. These assumptions ignore the fact that class difficulty evolves during learning (Sinha et al., 2022; Toneva et al., 2019). A class that initially seems difficult due to low representation may, over time, become easier to learn as the model adapts. Static correction cannot adapt to these dynamics, and often leads to overfitting on the minority class or wasting capacity by oversampling far beyond what is actually needed.
Figure 1: ART resampling
To this end, we propose ART, a resampling-based method that uses the class-wise macro F1-score as a simple heuristic for adjusting the data distribution. This is a novel approach that dynamically resamples data based on model performance. ART periodically monitors class-wise performance and updates the sampling distribution to emphasize underperforming classes. This class-level focus avoids the instability associated with instance-level difficulty estimation (Sinha et al., 2022) and treats class difficulty as dynamic rather than static throughout training. Figure1 shows how the sampling distribution of classes is changed based on the macro F1-score as training progresses. Using a class-wise metric as a resampling heuristic offers a simpler and more robust alternative to complex fixed heuristics. Unlike  (Sinha et al., 2022), which uses both resampling and a modified loss function, ART demonstrates that performance metric-based resampling alone outperforms existing methods. Also, ART utilizes a balance of performance metric-based sampling and the original dataset priors rather than completely ignoring the original data distribution. We demonstrate that ART consistently outperforms existing methods across classification tasks, achieving superior accuracy, recall, precision, and F1-scores. We also show that ART performs better than the baseline at varying levels of imbalance. Furthermore, we evaluate ART on classification tasks across tabular, image, and text datasets to demonstrate its modality-agnostic nature.
Unlike existing imbalance-handling approaches that rely on fixed resampling heuristics, instance-level difficulty estimation, or loss reweighting, ART introduces a metric-driven, class-level adaptive resampling strategy. By using periodically evaluated class-wise macro F1 scores as feedback, ART dynamically reallocates training focus based on evolving class difficulty rather than static class priors. This design makes ART model-agnostic, stable, and easy to integrate into standard training pipelines, while directly aligning training adaptation with commonly used imbalance evaluation metrics.
2 Related Work
@@
Resampling-based methods: A common approach to handle class imbalance is to rebalance the data distribution before or during training. Classic techniques include ROS, which simply duplicates minority-class examples, and RUS, which deletes majority-class examples. ROS and RUS are simple resampling methods and have been shown to be surprisingly competitive on some tasks. More sophisticated oversampling methods generate synthetic examples. These include methods like SMOTE (Chawla et al., 2002), which interpolates new minority examples in the feature space. Many SMOTE variants aim to focus sampling on hard or informative regions. For instance, borderline-SMOTE (Han et al., 2005) generates points near class boundaries, and ADASYN (He et al., 2008) adjusts sampling density based on local minority density. There are also some recent methods like safe-level-SMOTE (Bunkhumpornpat et al., 2009) and k-means–SMOTE (Douzas et al., 2018) that combine clustering or noise filtering with synthetic generation. In practice, oversampling can introduce noisy or redundant points, so hybrid methods that combine oversampling with cleaning have been proposed in (Batista et al., 2004b). For example, SMOTE-ENN and SMOTE-Tomek, as described in (Batista et al., 2004a), use edited nearest neighbors or Tomek-link removal to prune ambiguous samples after SMOTE. It has been shown that these hybrid methods yield well-defined class clusters and often outperform vanilla undersampling, especially when the minority class is minimal (Batista et al., 2004b, a).
@@
Unlike traditional resampling techniques that rely on fixed, data-level heuristics applied prior to training, ART dynamically adapts sampling probabilities during training based on evolving class-wise performance.
Loss-based methods: Rather than altering the data distribution, many techniques modify the learning algorithm or loss function. Classic cost-sensitive learning assigns a higher weight to the loss component of minority classes (Elkan, 2001a). One can weigh the loss inversely by class frequency, or use a pre-specified cost matrix (Elkan, 2001b). However, determining optimal costs can be difficult for complex models. Focal loss (Lin et al., 2018) addresses class imbalance in deep learning by changing the loss contribution of easy (well-classified) examples, thereby encouraging the model to focus on harder instances. Similarly, OHEM (Shrivastava et al., 2016) dynamically selects the examples with the highest loss within each mini-batch and focuses training on those “hard” samples rather than the full dataset. Although OHEM originated in computer vision, the idea of mining difficult instances can be applied in any domain.
While cost-sensitive and loss-reweighting methods modify the optimization objective, ART instead adjusts the data sampling process using external performance feedback, making it complementary to loss-based imbalance handling techniques.
@@
In contrast to margin-based approaches that enforce static class-dependent decision boundaries, ART adapts training emphasis dynamically according to observed class-level performance rather than predefined margins.
@@
Techniques like self-paced learning (Kumar et al., 2010) or focal loss adjust selection based on how confidently the model predicts each sample. Some recent works, like AdaSampling frameworks (Zhang et al., 2019), iteratively update example weights by model feedback, although many such methods focus on label-noise scenarios.
Unlike curriculum and self-paced learning approaches that rely on instance-level difficulty estimation or loss thresholds, ART operates at the class level and uses evaluation-aligned metrics, avoiding noisy instance-wise difficulty signals.
Margin-based sampling methods, like MSYN (Fan et al., 2011), explicitly choose instances near class boundaries to improve learning when the classifier has small margins on some samples. In general, instance-level adaptive sampling contrasts with class-level rebalancing because it can upweight or downweight individual examples even within the same class. Whereas instance-level difficulty methods focus on identifying and emphasizing individual hard samples, ART aggregates difficulty at the class level, resulting in more stable and interpretable adaptation in imbalanced settings.
Multi-modal Imbalanced Learning: In practice, imbalanced methods are applied across domains (vision, NLP, tabular data, etc.). Techniques like random sampling and SMOTE assume only vector inputs and have been used extensively in tabular and image datasets. Modern deep-learning loss reweighting (e.g., LDAM+DRW, focal loss) was popularized on vision tasks but extends naturally to NLP or other modalities. For instance, focal loss has been applied in text classification to handle rare classes, and cost-sensitive learning is common in medical diagnosis or fraud detection on tabular data. Class-balanced sampling and ensemble methods like Balanced Meta-Softmax (Ren et al., 2020) and LDAM+DRW have been evaluated on both image benchmarks and real-world text or tabular tasks. Moreover, some recent works explicitly benchmark multiple modalities: for example, Cui et al. (Cui et al., 2019) and Cao et al. (Cao et al., 2019) validated their class-balanced and margin-based losses on both image recognition and text classification tasks. Although prior dynamic strategies adjust weights or sampling schedules heuristically, ART uniquely employs class-wise macro F1 as an explicit feedback signal, directly aligning training adaptation with imbalance evaluation metrics.
Overall, existing deep imbalanced learning methods primarily focus on loss design or instance-level adaptation, whereas ART introduces a simple, model-agnostic class-level feedback mechanism that can be integrated into standard training pipelines.
3 Proposed Adaptive Resampling-based Training (ART) module
Figure  2 shows how training progresses under ART.
Figure 2: Overview of the training procedure in ART. The process begins with uniform initialization of class weights. Training proceeds in epochs, with periodic updates to class weights every bf epochs based on per-class macro F1-scores evaluated on a validation set.
@@ 3 Proposed Adaptive Resampling-based Training (ART) module
Algorithm 1 illustrates how these initial weights are blended with the empirical priors via convex interpolation to form the first sampling distribution. The classification model used in this study is a parameterized function, defined by:
@@
3.1 Overview
The aim of ART is to improve model performance on classification tasks by addressing variations in predictive performance across classes. ART accomplishes this by using an adaptive sampling mechanism informed by the classifier’s current performance, rather than relying solely on fixed empirical class frequencies. In contrast to conventional static resampling methods, ART periodically revises the sampling distribution by assessing and quantifying the classification difficulty of each class.
The algorithm for training and resampling under ART is explained in Algorithm 1. Classes identified as challenging, indicated by lower predictive performance, receive increased sampling priority. Consequently, ART systematically reallocates training resources towards improving performance on classes exhibiting greater classification complexity.
@@ 3.1 Overview
  1. 1.
    the blending constant, $c\in[0,1]$, controls the balance between sampling based on empirical class priors and the adaptive distribution derived from current class performance metrics;
  2. 2.
    the boosting frequency, $bf\in\mathbb{Z}^{+}$, specifies the interval, measured in epochs, at which the sampling distribution is updated.
These hyperparameters influence how ART adjusts its sampling strategy over the course of training and affects both the frequency and intensity of its response to class-level learning difficulty changes.
@@
3.2 Theoretical Motivation and Optimization Perspective of ART
ART can be interpreted as a dynamic reweighting scheme that operates directly on the data distribution rather than on the loss function. From an optimization perspective, standard empirical risk minimization (ERM) on imbalanced data minimizes a weighted sum of per-class risks, where the weights are implicitly determined by the empirical class priors. As a result, majority classes dominate the optimization objective, while minority classes contribute weakly to the gradient updates, even when their predictive performance remains poor.
Formally, training on the original dataset corresponds to minimizing the following empirical objective:
$\mathcal{R}(\theta)=\sum_{i=0}^{K-1}\Pi_{i}\,\mathbb{E}_{(x,y)\sim D_{i}}\left[\mathcal{L}(f_{\theta}(x),y)\right],$
where $\Pi_{i}$ denotes the empirical prior of class $i$. This objective is static: The contribution of each class to the optimization process is fixed throughout training, regardless of how class-wise performance evolves. This often results in situations where some classes remain under-optimized even after convergence, especially in severely imbalanced settings.
ART generalizes this by replacing the fixed class prior $\Pi_{i}$ with a time-varying sampling distribution $p_{i}^{(t)}$, which is updated based on the model’s observed class-wise performance.
At the training step $t$, ART approximately optimizes the given objective:
$\mathcal{R}^{(t)}(\theta)=\sum_{i=0}^{K-1}p_{i}^{(t)}\,\mathbb{E}_{(x,y)\sim D_{i}}\left[\mathcal{L}(f_{\theta}(x),y)\right].$
This highlights that ART performs dynamic ERM, adjusting the relative importance of each class in response to the model’s current deficiencies. Unlike loss-based reweighting methods, which modify gradient magnitudes directly, ART achieves this reweighting indirectly by altering the frequency with which samples from each class are observed during training.
The update rule for $p_{i}^{(t)}$ is driven by class-wise macro F1-scores evaluated on a held-out validation set. The macro F1-score provides a balanced measure of class-level predictive quality by jointly capturing both precision and recall. Importantly, macro F1 is insensitive to class frequency, ensuring that poor performance on a minority class is not obscured by its limited representation. A low macro F1-score for a class indicates that the model is
  • 1.
    either failing to correctly identify instances of that class (low recall),
  • 2.
    misclassifying many predictions into that class (low precision), or
  • 3.
    both these cases,
making it a reliable indicator of class-level learning difficulty.
By defining the difficulty score as $s_{i}=1-f_{i}$, ART establishes a monotonic relationship between performance and sampling priority: classes with lower predictive performance receive higher sampling probability in subsequent training phases. This design aligns the resampling mechanism with the evaluation objective, ensuring that optimization effort is concentrated on the classes that most negatively impact the macro-level performance metric. From this perspective, ART can be viewed as a feedback-driven training procedure, in which validation-time performance signals guide future data exposure.
3.3 Performance-based Sampling
@@
These computed difficulty scores reflect the inverse performance relationship and hence, emphasize classes with poor predictive performance. To create a valid probability distribution from these difficulty scores, we normalize them across all classes, resulting in performance-based sampling weights given by:
@@
3.4 Blending with Class Priors